# Pushing the Envelope of LLM Inference on AI-PC

### v1

@@ Abstract
The advent of ultra-low-bit LLM models (1/1.58/2-bit), which match the perplexity and end-task performance of their full-precision counterparts using the same model size, is ushering in a new era of LLM inference for resource-constrained environments such as edge devices and AI PCs. While these quantization advances promise models that are more cost-effective in terms of latency, memory, throughput, and energy consumption, the computational efficiency of state-of-the-art (SOTA) inference runtimes (e.g., bitnet.cpp) used to deploy them remains underexplored. In this work, we take a bottom-up approach: we first design and implement 1-bit and 2-bit microkernels optimized for modern CPUs, achieving peak computational efficiency across a variety of CPU platforms. We integrate these microkernels into a state-of-the-art LLM inference framework, namely PyTorch-TPP, and present end-to-end inference results with 2-bit models that outperform the current SOTA runtime bitnet.cpp by up to 2.2, and deliver up to 7 speedup compared to the 16-bit model inference. Our optimized runtime advances the state of LLM inference on AI PCs and edge devices, paving the way for efficient deployment of ultra-low-bit LLM models.
@@ I Introduction
Recent advances in Quantization-Aware Training (QAT) LLM methods have produced ultra-low bit (1-bit, 1.58-bit and 2-bit) quantized models that match the perplexity and end-task performance of their full-precision counterparts using the same model size [1, 2]. Despite these advancements that yield ultra-low bit LLM models with the potential to be cost-effective in terms of latency, memory, throughput, and energy consumption, the available runtimes capable of serving inference with such models are limited. For example, the state-of-the-art inference runtime bitnet.cpp is specifically designed for ternary LLMs, with a relatively narrow range of applicable model architectures [3]. Additionally, although the bitnet.cpp runtime improves upon alternatives (e.g. llama.cpp) being up to 6 faster by introducing a specialized mixed-precision matrix multiplication (mpGEMM) library to facilitate sub-2-bits model inference, its efficiency is not evaluated in terms of optimality or roofline performance. Our preliminary performance analysis showed that 2-bit inference with bitnet.cpp was even slower that state-of-the-art 4-bit inference on CPUs (e.g. see in Figures 9, 8, 10 the orange and the magenta bars), highlighting that the bitnet.cpp runtime is far from optimal.
@@
We designed and implemented 1-bit and 2-bit GEMM kernels tailored for CPUs with AVX2 ISA. These kernels leverage novel blocked and interleaved weight tensor layouts to facilitate the up-convert process during inference. We devised performance models that characterize the effectiveness and limitations of these kernels.
We composed these new microkernels into multi-threaded GEMM routines targeting hybrid-core architectures. We benchmarked the new routines on a set of matrices and platforms, showing that they can achieve the attainable memory bandwidth, effectively accelerating the bandwidth-bound matrix-vector multiplication operations proportionally to the bit-width of the weight’s datatypes.
We integrated these ultra-low bit GEMM kernels in our SOTA end-to-end inference pipeline, assessed its performance on a range of models and platforms, and observed speedups up to 7 over the 16-bit inference.
We benchmarked our runtime against bitnet.cpp and show that our proposed solution is up to 2.2 faster. Finally, we compare our CPU performance with the optimized GPU 2-bit inference within bitnet.cpp, which includes kernels specifically tailored for the A100 platform. Our performance on contemporary x86 AI-PC CPUs is within 2.3-3 of the A100 GPU, which has 17-20 more bandwidth than the CPU platforms. Thus, our work pushed the envelope of LLM inference on AI-PCs and demonstrates that ultra-low-bit inference on CPUs can approach GPU-level performance.
@@ III-A Overview and background
In this section we detail the ultra-low bit GEMM microkernels, that take as input an 1-bit (int1) or 2-bit (int2) weight tensor, multiply it with the input activation tensor and produce the output activations. We refer to this technique as ”up-convert and compute”. We focus on modern CPUs with AVX2 instructions that can be found on contemporary AI-PC devices. As described in Section I, we target the use-cases where the GEMMs illustrate algorithmically very low arithmetic intensity and are typically bandwidth-bound (i.e. single-batch inference). We implemented all the GEMM microkernels using the libxsmm library, a Just-In-Time (JIT) GEMM library that delivers SOTA GEMM performance on CPUs [23, 4]. Last but not least, our multi-threaded GEMMs use these JIT-ed microkernels as building blocks, and for parallelization we leverage PARLOOPER and its hybrid scheduling capabilities to fully utilize all performance and efficiency cores [18].
III-B Interleaved Tensor Layout for 2-bit Weights
@@
To address this, we quantize the input activations to int8, allowing us to leverage the int8 VNNI FMAs available on modern AVX2 CPUs. These int8 VNNI FMAs offer the throughput of FP32 FMAs. The int8 VNNI FMA requires the input vector operands to pack together four 8-bit int8 values from the logical contraction dimension of the GEMM (which we denote as throughout the rest of the paper), and this is the so-called VNNI4 layout. Naively packing together four 2-bit values in the innermost dimension of the weight tensor to effectively form a VNNI4 layout, requires during runtime an excessive number of shifts and shuffles to (i) unpack the int2 values to int8, and (ii) to re-pack the int8 values back to VNNI4 layout (with a two-level shuffle network) for the upcoming int8-VNNI FMAs.
@@
To alleviate this unpacking bottleneck in the case of 2-bit weights, we introduce a new tensor layout called VNNI4-interleaved, illustrated in Figure 1. In this layout, a logical tensor is transformed into an interleaved 3D tensor by folding in the innermost dimension entries from the logical dimension that are iso-modulo 8. The new tensor layout reflects a 3D tensor whereas the innermost folded dimension corresponds to the originally-indexed entries with the same modulo 8 index. In this way, we pack a 2-bit subtensor in a 32-bit value (see doted sub-tensor in top of Figure 1), and with a 256-bit vector-load we can read 128 2-bit entries, which effectively form an tensor.
@@
Using this layout we can simply perform table-lookups that utilize the 2-bit entries as indices and translate them to int8 values that effectively form within a 256-bit vector an int8 subtensor in VNNI4 format. This int8 vector can be readily used as input to int8 VNNI FMA. In the example of unpack instruction sequence (next to blue arrow in Figure 1), with the byte-shuffle highlighted in blue vpshufb ymm3, ymm_lut01, ymm3 we consider the 32 2-bit indices highlighted in blue from the original 2-bit packed vector. The ymm_lut01 considers only the bits 0-1 of each byte in the vector and maps these indices to the corresponding int8 values. In an analogous way, with the byte-shuffle highlighted in orange vpshufb ymm4, ymm_lut23, ymm3 we consider the 32 2-bit indices with orange color from the original 2-bit packed vector. The ymm_lut23 considers only bits 2-3 of each byte in the vector and maps these indices to the proper int8 values. The shift instruction vpsrld ymm5, ymm3, 0x4 shifts the 4 higher bits 4-7 within each byte to the 4 lower positions such that with 2 more shuffles (highlighted with green and magenta) we can unpack the corresponding int2 values to the respective int8 values. The two logical AND vector operations (vpand) in the unpack instruction sequence merely mask out the highest 4 bits within each byte to ensure that the byte-shuffles have the desired effect. After this instruction sequence with 1 logical shift + 2 logical AND + 4 byte-shuffles we get as output 4 256-bit vectors, each holding an int8 subtensor which is in VNNI4 layout.
@@
By leveraging the interleaved format described in section III-B we can design a mixed precision GEMM microkernel that uses 2-bit weights, 8-bit activations and produces 32-bit outputs (see Figure 2). First, with a full 256-bit vector load we load 128 int2 values in layout in vector register ymm3. Then with the up-convert sequence described in the previous section we unpack these 128 int2 values to 128 int8 values occupying 4 vector registers ymm3 - ymm6. With a 32-bit broadcast vpbroadcastd we broadcast in ymm7 four 8-bit values from the input activations that correspond to 4 elements from the logical dimension that should be contracted. Finally, with four vnni-int8 FMAs (vpdpbssd) we perform the corresponding multiply-add operations among vectors ymm3 - ymm6 and vector ymm7 and we accumulate the partial products in the output 32-bit accumulators ymm12 - ymm15. This microkernel effectively multiplies a logical subtensor with a vector and produces an output with 32-bit integer precision. Using this microkernel as a building block, we can compose GEMM and BRGEMM microkernels with larger , and values [23, 24].
III-D 1-bit GEMM microkernel
@@
For the 1-bit GEMM microkernel we use a conventional VNNI4 layout, where 32 int1 values are packed in an subtensor. Figure 3 illustrates a microkernel that multiplies 1-bit matrix in VNNI4 fromat with size with an 8-bit vector . First with a 32-bit broadcast vpbroadcastd ymm3,dword ptr[A] we broadcast a logical subtensor of to within each 32-bit chunk of vector ymm3. With an in-lane byte shuffle vpshufb ymm3, ymm3, ymm_perm we move the byte containing the -th bit of the loaded 32-bits to the -th byte within the 256-bit vector ymm3. Then with a logical AND operation vpand ymm3, ymm1, ymm3 we isolate within each byte the bit residing in the -th position. After this step, we use a compare instruction vpcmpeqb ymm3, ymm1, ymm3 to use the previous isolated bits and generate a mask with byte values 0x00 or 0xff depending on the isolated values of the bits. Last, by using the generated mask we blend with the instruction vpblendvb ymm3, ymm2, ymm3, ymm3 the int8 values +1/-1 to the expanded vector ymm3. With such an instruction sequence we upconvert the logical 1-bit subtensor of to an 8-bit subtensor in ymm3 that has the value +1 if the original bit in is 0 and the value -1 if the original bit in is 1 (this is the desired weight encoding in our 1-bit weights [2]). With a 32-bit broadcast vpbroadcastd we broadcast in ymm7 four 8-bit values from the input activations that correspond to 4 elements from the logical dimension that should be contracted. Finally, with a vnni-int8 FMAs (vpdpbssd ymm12, ymm3, ymm7) we perform the corresponding multiply-add operations among vectors ymm3 and vector ymm7 and we accumulate the partial product in the output 32-bit accumulator ymm12. We repeat the same up-convert and FMA sequence for the rest 96 entries of matrix (where we reuse the input vector ) and we conclude the int1 GEMM microkernel. With this microkernel as a building block, we compose GEMM and BRGEMM microkernels with larger , and values [24].
III-E Performance modeling

### v2

@@ Abstract
The advent of ultra-low-bit LLM models (1/1.58/2-bit), which match the perplexity and end-task performance of their full-precision counterparts using the same model size, is ushering in a new era of LLM inference for resource-constrained environments such as edge devices and AI PCs. While these quantization advances promise models that are more cost-effective in terms of latency, memory, throughput, and energy consumption, the computational efficiency of state-of-the-art (SOTA) inference runtimes (e.g., bitnet.cpp) used to deploy them remains underexplored. In this work, we take a bottom-up approach: we first design and implement 1-bit and 2-bit microkernels optimized for modern CPUs, achieving peak computational efficiency across a variety of CPU platforms. We integrate these microkernels into a state-of-the-art LLM inference framework, namely PyTorch-TPP, and present end-to-end inference results with 2-bit models that outperform the current SOTA runtime bitnet.cpp by up to 2.2, and deliver up to 7 speedup compared to the 16-bit model inference. We then extend this work to Intel GPUs where we design and implement mixed precision, 2-bit GEMM kernels, and show their performance to be close to optimal. We integrated our optimized Xe2 kernels in the vLLM framework as a quantization plugin and evaluated end-to-end LLM inference results for a range of LLM models and Xe2 GPUs. Depending on the model and platform, we see a 4 - 8 reduction in GEMM time compared to the BF16 case, and we get up to 6.3 speedup in end-to-end latency compared to the BF16 execution. Our optimized runtime advances the state of LLM inference on AI PCs and Intel Xe GPUs, paving the way for efficient deployment of ultra-low-bit LLM models.
@@ I Introduction
Recent advances in Quantization-Aware Training (QAT) LLM methods have produced ultra-low bit (1-bit, 1.58-bit and 2-bit) quantized models that match the perplexity and end-task performance of their full-precision counterparts using the same model size [1, 2]. Despite these advancements that yield ultra-low bit LLM models with the potential to be cost-effective in terms of latency, memory, throughput, and energy consumption, the available runtimes capable of serving inference with such models are limited. For example, the state-of-the-art inference runtime bitnet.cpp is specifically designed for ternary LLMs, with a relatively narrow range of applicable model architectures [3]. Additionally, although the bitnet.cpp runtime improves upon alternatives (e.g. llama.cpp) being up to 6 faster by introducing a specialized mixed-precision matrix multiplication (mpGEMM) library to facilitate sub-2-bits model inference, its efficiency is not evaluated in terms of optimality or roofline performance. Our preliminary performance analysis showed that 2-bit inference with bitnet.cpp was even slower than state-of-the-art 4-bit inference on CPUs (e.g. see in Figures 13, 12, 14 the orange and the magenta bars), highlighting that the bitnet.cpp runtime is far from optimal.
@@
Then we extend this work to Intel Xe2 GPUs where we design and implement mixed precision, 2-bit GEMM kernels. While contemporary Xe2 Intel GPUs support natively int2int8int32 GEMM computations using hardware acceleration, the SOTA ultra-low bit inference (e.g. 2-bit weights) requires high-precision (e.g. BF16) input/output activations, i.e. int2BF16BF16 GEMM. Quantizing the input activations from BF16 to int8 before the int2int8int32 GEMM operation, and dequantizing the int32 GEMM output to BF16 output activations via separate microkernels out of L1/L2 cache adds minimal overhead for the CPU use-case. However, on GPUs such a decoupled execution of activation-quantization low-precision GEMM microkernel output-dequantization incurs multiple kernel launches and passes over the data. Therefore, we developed mixed precision int2 BF16BF16 GEMM Xe2 kernels that fuse the input/output activation quantization/dequantization within the GEMM kernel to minimize the corresponding overheads. We integrated our optimized Xe2 kernels in the vLLM [5] framework as a quantization plugin and evaluated end-to-end LLM inference results for a range of LLM models and Intel Xe2 GPUs. Depending on the model and platform, we see a 4-8 reduction in GEMM time compared to the BF16 case, and we get up to 6.3 speedup in end-to-end latency compared to the BF16 execution.
@@
We designed and implemented 1-bit and 2-bit GEMM kernels tailored for CPUs with AVX2 ISA. These kernels leverage novel blocked and interleaved weight tensor layouts to facilitate the up-convert process during inference. We devised performance models that characterize the effectiveness and limitations of these kernels. We composed these new CPU microkernels into multi-threaded GEMM routines targeting hybrid-core architectures. Also, we benchmarked the new routines on a set of matrices and platforms, showing that they can achieve the attainable memory bandwidth, effectively accelerating the bandwidth-bound matrix-vector multiplication operations proportionally to the bit-width of the weight’s datatypes.
We designed and implemented mixed precision GEMM kernels targeting Intel GPUs with Xe2 cores. These kernels fuse the quantization/dequantization of the input/output activations and leverage the hardware accelerated int2int8 Dot Product Accumulate Systolic (DPAS) instructions. We benchmarked the mixed precision GEMM kernels under different scenarios and show they can achieve bandwidth close to the platform’s limit, accelerating matrix-vector multiplication operations proportionally to the bit-width of the weight’s datatypes. We also achieve compute throughput close to the ideal int2int8 DPAS throughput, highlighting the efficacy of our fused quantization/dequantization, and these Xe2 GEMM kernels can be further used to accelerate compute-bound use-cases in LLM inference (e.g. prefill phase, larger batch-size inference).
We integrated these ultra-low bit GEMM kernels in SOTA end-to-end inference pipelines (PyTorch-TPP for CPU, vLLM for GPU), assessed their performance on a range of models and platforms, and observed speedups up to 7 over the 16-bit inference. We benchmarked our runtime against bitnet.cpp and show that our proposed solution is up to 2.2 faster.
We compare our CPU performance with the optimized GPU 2-bit inference within bitnet.cpp, which includes kernels specifically tailored for the NVIDIA A100 platform. Our performance on contemporary x86 AI-PC CPUs is within 2.3-3 of the A100 GPU, which has 17-20 more bandwidth than the CPU platforms. On a discrete Intel Arc B580 Xe2 GPU we achieve a speedup of 1.5 over the 2-bit inference on A100 for a same-sized 2 Billion parameter model, despite A100 having 4 more bandwidth than B580. Thus, our work pushes the envelope of LLM inference on AI-PCs and Intel Xe GPUs, and demonstrates that ultra-low-bit inference even on CPUs and discrete client GPUs can approach high-end GPU-level performance.
@@ III-A Overview and background
In this section we detail the ultra-low bit GEMM microkernels, that take as input an 1-bit (int1) or 2-bit (int2) weight tensor, multiply it with the input activation tensor and produce the output activations. We refer to this technique as ”up-convert and compute”. We focus on modern CPUs with AVX2 instructions that can be found on contemporary AI-PC devices, and also on modern Intel GPUs with Xe2 cores that natively support int2int8 operations. As described in Section I, we target the use-cases where the GEMMs illustrate algorithmically very low arithmetic intensity and are typically bandwidth-bound (i.e. single-batch inference). For the CPU case, we implemented all the GEMM microkernels using the libxsmm library, a Just-In-Time (JIT) GEMM library that delivers SOTA GEMM performance on CPUs [24, 4]. Our multi-threaded GEMMs for CPUs use these JIT-ed microkernels as building blocks, and for parallelization we leverage PARLOOPER and its hybrid scheduling capabilities to fully utilize all performance and efficiency cores [19]. For the Xe2 GPUs, we implement the corresponding low-bit microkernels using the XeTLA framework: a collection of SYCL/ESIMD templates that enable high-performance computations on Intel Xe GPU architectures [25].
III-B Interleaved Tensor Layout for 2-bit Weights on CPU
@@
To address this, we quantize the input activations to int8, allowing us to leverage the int8 VNNI FMAs available on modern AVX2 CPUs that offer the throughput of FP32 FMAs. The input activation quantization, along with the required output activation dequantization can be implemented as separate operations before/after the quantized GEMM operations with minimal overhead on the CPUs. More specifically, for memory-bound matrix-vector multiplication scenarios the input and output activations are merely vectors with small size compared to the weight tensors, and as such their quantization and dequantization incur negligible overhead. The int8 VNNI FMA requires the input vector operands to pack together four 8-bit int8 values from the logical contraction dimension of the GEMM (which we denote as throughout the rest of the paper), and this is the so-called VNNI4 layout. Naively packing together four 2-bit values in the innermost dimension of the weight tensor to effectively form a VNNI4 layout, requires during runtime an excessive number of shifts and shuffles to (i) unpack the int2 values to int8, and (ii) to re-pack the int8 values back to VNNI4 layout (with a two-level shuffle network) for the upcoming int8-VNNI FMAs.
@@
To alleviate this unpacking bottleneck in the case of 2-bit weights, we introduce a new tensor layout called VNNI4-interleaved, illustrated in Figure 1-Left. In this layout, a logical tensor is transformed into an interleaved 3D tensor by folding in the innermost dimension entries from the logical dimension that are iso-modulo 8. The new tensor layout reflects a 3D tensor whereas the innermost folded dimension corresponds to the originally-indexed entries with the same modulo 8 index. In this way, we pack a 2-bit subtensor in a 32-bit value (see dotted sub-tensor in top of Figure 1-Left), and with a 256-bit vector-load we can read 128 2-bit entries, which effectively form an tensor.
@@
Using this layout we can simply perform table-lookups that utilize the 2-bit entries as indices and translate them to int8 values that effectively form within a 256-bit vector an int8 subtensor in VNNI4 format. This int8 vector can be readily used as input to int8 VNNI FMA. In the example of unpack instruction sequence (between the arrows in Figure 1-Left), with the byte-shuffle highlighted in blue vpshufb ymm3, ymm_lut01, ymm3 we consider the 32 2-bit indices highlighted in blue from the original 2-bit packed vector. The ymm_lut01 considers only the bits 0-1 of each byte in the vector and maps these indices to the corresponding int8 values. In an analogous way, with the byte-shuffle highlighted in orange vpshufb ymm4, ymm_lut23, ymm3 we consider the 32 2-bit indices with orange color from the original 2-bit packed vector. The ymm_lut23 considers only bits 2-3 of each byte in the vector and maps these indices to the proper int8 values. The shift instruction vpsrld ymm5, ymm3, 0x4 shifts the 4 higher bits 4-7 within each byte to the 4 lower positions such that with 2 more shuffles (highlighted with green and magenta) we can unpack the corresponding int2 values to the respective int8 values. The two logical AND vector operations (vpand) in the unpack instruction sequence merely mask out the highest 4 bits within each byte to ensure that the byte-shuffles have the desired effect. After this instruction sequence with 1 logical shift + 2 logical AND + 4 byte-shuffles we get as output 4 256-bit vectors, each holding an int8 subtensor which is in VNNI4 layout.
@@
By leveraging the interleaved format described in section III-B we can design a mixed precision GEMM microkernel that uses 2-bit weights, 8-bit activations and produces 32-bit outputs (see Figure 1-Right). First, with a full 256-bit vector load we load 128 int2 values in layout in vector register ymm3. Then with the up-convert sequence described in the previous section we unpack these 128 int2 values to 128 int8 values occupying 4 vector registers ymm3 - ymm6. With a 32-bit broadcast vpbroadcastd we broadcast in ymm7 four 8-bit values from the input activations that correspond to 4 elements from the logical dimension that should be contracted. Finally, with four vnni-int8 FMAs (vpdpbssd) we perform the corresponding multiply-add operations among vectors ymm3 - ymm6 and vector ymm7 and we accumulate the partial products in the output 32-bit accumulators ymm12 - ymm15. This microkernel effectively multiplies a logical subtensor with a vector and produces an output with 32-bit integer precision. Using this microkernel as a building block, we can compose GEMM and BRGEMM microkernels with larger , and values [24, 26].
III-D 1-bit GEMM microkernel for CPU
@@
For the 1-bit GEMM microkernel we use a conventional VNNI4 layout, where 32 int1 values are packed in an subtensor. Figure 2 illustrates a microkernel that multiplies 1-bit matrix in VNNI4 format with size with an 8-bit vector . First with a 32-bit broadcast vpbroadcastd ymm3,dword ptr[A] we broadcast a logical subtensor of to within each 32-bit chunk of vector ymm3. With an in-lane byte shuffle vpshufb ymm3, ymm3, ymm_perm we move the byte containing the -th bit of the loaded 32-bits to the -th byte within the 256-bit vector ymm3. Then with a logical AND operation vpand ymm3, ymm1, ymm3 we isolate within each byte the bit residing in the -th position. After this step, we use a compare instruction vpcmpeqb ymm3, ymm1, ymm3 to use the previous isolated bits and generate a mask with byte values 0x00 or 0xff depending on the isolated values of the bits. Last, by using the generated mask we blend with the instruction vpblendvb ymm3, ymm2, ymm3, ymm3 the int8 values +1/-1 to the expanded vector ymm3. With such an instruction sequence we upconvert the logical 1-bit subtensor of to an 8-bit subtensor in ymm3 that has the value +1 if the original bit in is 0 and the value -1 if the original bit in is 1 (this is the desired weight encoding in our 1-bit weights [2]). With a 32-bit broadcast vpbroadcastd we broadcast in ymm7 four 8-bit values from the input activations that correspond to 4 elements from the logical dimension that should be contracted. Finally, with a vnni-int8 FMAs (vpdpbssd ymm12, ymm3, ymm7) we perform the corresponding multiply-add operations among vectors ymm3 and vector ymm7 and we accumulate the partial product in the output 32-bit accumulator ymm12. We repeat the same up-convert and FMA sequence for the remaining 96 entries of matrix (where we reuse the input vector ) and we conclude the int1 GEMM microkernel. With this microkernel as a building block, we compose GEMM and BRGEMM microkernels with larger , and values [26].
III-E 2-bit GEMM kernels for Intel Xe2 GPU
Contemporary Xe2 Intel GPUs support natively int2int8int32 GEMM computations using hardware-accelerated Dot Product Accumulate Systolic (DPAS) instructions. Therefore, for the Xe2 int2 kernels we can natively use these instructions without the need to upconvert the weights to int8 before performing the necessary multiply-add operations with the int8 activations (cf. with the CPU up-convert& compute methodology). Nevertheless, the SOTA ultra-low bit inference (e.g. 2-bit weights) requires high-precision (e.g. BF16) input/output activations, i.e. the involved GEMM precision is int2BF16BF16. For the CPU low-precision GEMM, quantizing the input activations from BF16 to int8 before the int2int8int32 GEMM operation, and dequantizing the int32 GEMM output to BF16 output activations via separate microkernels out of L1/L2 cache adds minimal overhead. More specifically, for memory-bound matrix-vector multiplication scenarios the input and output activations are merely vectors with small size compared to the weight tensors, therefore their quantization and dequantization incur negligible overhead since these operations may run out of L1/L2 cache. However, on GPUs such a decoupled execution (i.e. multiple kernels) of activation-quantization low-precision GEMM microkernel output-dequantization incurs multiple kernel launches and passes over the data.
To mitigate the shortcomings of multiple kernel launches, we developed mixed precision int2 BF16BF16 GEMM Xe2 kernels that fuse the input/output activation quantization/dequantization within the GEMM kernel (see Figure 3). Using conventional SYCL terminology [27], we split the output matrix into tiles with size and each workgroup will calculate a sub-matrix , represented by the yellow box in output C (see Left part of Figure 3). Subsequently, this sub-matrix will be continuously divided into multiple tiles, with dimensions . These tiles will then be assigned to subgroups. Finally, the corresponding subgroup GEMM microkernel and the involved tile operations will be mapped to the actual Xe2 hardware instructions, such as 2D-load and DPAS instructions. The Xe2 int2int8int32 DPAS instructions for Xe2 require the int2 operand to be formatted in VNNI16 layout: pack together 16 int2 values from the logical contraction dimension of the GEMM. Since we are performing inference we can merely pre-format the int2 weights to such VNNI16 layout without any overhead during runtime.
Figure 3 also illustrates the fusion of the input activation (matrix ) quantization and the output quantization. When loading the matrix from memory in registers, the precision is BF16 (see dark blue box). In order to quantize the BF16 entries given some scales (see blue box in the GEMM microkernel assembly), we first multiply the entries with the inverse of the scales, then we convert the result to int8 (with saturation) and these int8 values (light-blue box) can be readily used in the DPAS instructions (see peach-colored box in microkernel assembly). By carefully selecting the shapes of the tiles we can re-use the quantized entries and amortize the corresponding overhead (for the rectangular tiles of Figure 3 we re-use each quantized register 8 times). We iterate the same process across all the tiles in the dimension. Once the full accumulation has been completed (either at the full granularity or at some user-defined granularity), we need to dequantize the int32 output registers forming a logical sub-tensor. First we convert the int32 values to float, then we multiply these float values with the dequantization scales and finally we convert the result to BF16 than can be stored to the output in memory e.g. via 2D-store Xe2 operations (see purple box in microkernel assembly). Also any additional operations (e.g. bias addition, post-GEMM activation functions) can be fused at this point where the entries are still in-registers before storing them to memory. Our Xe2 GEMM implementation also supports -splitting algorithms that extract parallelism across the dimension for smaller GEMM shapes [25].
With respect to the calculation of the , we have to compute the absolute maximum per quantization granularity (e.g. the absolute max of entries or some other user-defined granularity). For this abs-max operation we implemented an optimized SYCL reduction kernel that can be either performed before the GEMM as a separate kernel or it can be fused at the work-group level, and the calculated abs-max values are stored in fast Shared Local Memory (SLM) available on Xe2. Fusing the abs-max reduction in the GEMM kernel avoids one kernel-launch overhead at the expense of some redundant computations (since multiple work-groups might be performing the same reduction in case they share the same slabs of ). We expose this option of fusing/non-fusing scale calculation as a tuning knob in our GEMM implementation. Also, the C dequantization scales mentioned earlier are merely a product of the corresponding and the that are part of the input int2 weight matrix.
Last but not least, in order to extract the maximum performance of our int2 Xe2 GEMM kernels for a variety of GEMM shapes and Xe2 platforms, we developed an auto-tuning framework within XeTLA that searches for optimal kernel parameter configurations, e.g. fusing/non-fusing scale calculation, optimal , , , tile sizes, using/not-using -splitting algorithms etc.
III-F Performance modeling & limitations of the CPU kernels